When given the domain {-2,-1,0,2,5}, solve the equation y= 3x+2.
Answer by sdmmadam@yahoo.com(530)

x belongs to the set: {-2,-1,0,2,5}
y= 3x+2. ----(1)
x=-2 in (1), y = 3X(-2)+2 =-6+2 = -4
x=-1 in (1), y = 3X(-1)+2 =-3+2 = -3+2 =-1
x= 0 in (1), y = 3X(0)+2 = 0+2 = 2
x=2 in (1), y = 3X(2)+2 = 6+2 = 8
x=5 in (1), y = 3X(5)+2 = 17+2 = 19
Therefore the solution set is given by
{-4,-1,2,8,17}